# Variational Quantum Framework for Partial Differential Equation Constrained Optimization

### v1

@@ Abstract
We present a novel variational quantum framework for partial differential equation (PDE) constrained design optimization problems. Such problems arise in simulation based design in many scientific and engineering domains. For instance in aerodynamic design, the PDE constraints are the conservation laws such as momentum, mass and energy balance, the design variables are vehicle shape parameters and material properties, and the objective could be to minimize the effect of transient heat loads on the vehicle or to maximize the lift. The proposed framework utilizes the variational quantum linear system (VQLS) algorithm and a black box optimizer as its two main building blocks. VQLS is used to solve the linear system, arising from the discretization of the PDE constraints for given design parameters, and evaluate the design cost/objective function. The black box optimizer is used to select next set of parameter values based on this evaluated cost, leading to nested bi-level optimization structure within a hybrid classical-quantum setting. We present detailed complexity analysis to highlight the potential advantages of our proposed framework over classical techniques. We implement our framework using the PennyLane library, apply it to solve a prototypical heat transfer optimization problem, and present simulation results using Bayesian optimization as the black box optimizer.
@@ 1 Introduction
We present a novel variational quantum framework for optimization problems constrained by partial differential equations (PDE). Variational quantum algorithms (VQAs) are hybrid classical-quantum algorithms that have emerged as promising candidates to optimally utilize today’s Noisy Intermediate Scale Quantum (NISQ) devices. VQAs use a quantum subroutine to evaluate the cost/ energy function using a shallow parameterized quantum circuit. They provide a versatile framework to tackle a wide variety of problems. As such, they have been successfully used in many applications such as finding ground and excited states of Hamiltonians [1, 2, 3, 4], dynamical quantum simulation [5, 6], combinatorial optimization [7, 8], solving linear systems [9, 10, 11], integer factorization [12], principal component analysis [13, 14, 15] and quantum machine learning [16].
@@
In particular, we are interested in the application of the variational quantum linear solver (VQLS) [9], a type of VQA for solving ordinary differential equations (ODEs) and partial differential equations (PDEs) [17, 18, 19, 20]. In this approach the ODEs/PDEs are discretized in space/time and embedded into a system of linear equations. VQLS is then applied to the linear system to obtain the normalized solution and compute observables of interest. While there are no theoretical guarantees associated with VQLS, its query complexity is empirically found to scale polylogarithmically with the size of the linear system [9] and thus could be beneficial over classical linear solvers that have a polynomial scaling.
In this work, we extend the VQLS based ODE/PDE solution framework for PDE constrained design optimization problems. Such problems arise in simulation based design, for instance in aerodynamic design using computational fluid dynamics (CFD) simulations. In this context the PDE constraints are the conservation laws such as momentum, mass and energy balance, the design variables are aerodynamic shape parameters and material properties, and the objective is to minimize heat transfer or to maximize lift/minimize drag. Since CFD simulations are computationally intensive, using them within the design/optimization loop can be prohibitively expensive. To address this challenge we propose a variational quantum framework where the simulation and design problem can be integrated and solved simultaneously. The main contributions of this work are as follows:
@@
  • We present detailed complexity analysis for VQLS based solution of linear ODEs under explicit and implicit Euler discretization schemes, and use that to assess potential advantage of our BVQPCO framework over classical techniques.
  • We implement our framework using the PennyLane library, apply it to solve a prototypical heat transfer optimization problem, and present simulation results using Bayesian optimization (BO) as the black box optimizer.
In Sec. 3 we motivate the general structure of PDE constrained design optimization problem through a detailed mathematical formulation of the 1D heat transfer optimization problem. Sec. 4.1 and 4.2 provide an overview of VQLS and BO, respectively. We describe our BVQPCO framework and outline a pseudo-code for its implementation in Sec. 5. We present the main results from the complexity analysis of the BVQPCO framework in Sec. 6; detailed proofs are provided in the Appendix. In  Sec. 7 we illustrate the BVQPCO framework on the heat transfer optimization problem and provide details on the implementation using the PennyLane library. We finally conclude in Sec. 8 with avenues for future work.
@@ 2 Notation
We will use standard braket notation, i.e. $|\mathbf{\psi}\rangle$ and $\langle\mathbf{\psi}|$ in representing the quantum states. The inner product between two quantum states $|\mathbf{\psi}\rangle$ and $|\mathbf{\phi}\rangle$ will be denoted by $\langle\mathbf{\psi}|\mathbf{\phi}\rangle$. For a vector $\mathbf{x}$, we denote by $|\mathbf{x}\rangle=\frac{\mathbf{x}}{\|\mathbf{x}\|}$ as the corresponding quantum state. The trace norm $\rho(\mathbf{\psi},\mathbf{\phi})$ between two pure states $|\mathbf{\psi}\rangle$ and $|\mathbf{\phi}\rangle$ is defined as
@@ 3 PDE Constrained Optimization
One popular classical approach to solve PDE constrained optimization is via gradient based methods [22, 23]. For example, the gradient of the cost function $C_{d}$ w.r.t to $\mathbf{p}$ can be obtained via finite differencing. However, that requires solving the PDE (8) at least $N_{p}+1$ times to approximate the derivative of solution $\mathbf{u}$ w.r.t. the parameters. A more efficient approach is via adjoint formulation, where one needs to solve the PDE (8) only one time and an additional adjoint equation to compute the required gradient [22, 23]. However, given that this gradient based approach requires forward simulation(s) for each gradient step, it can be computationally intensive. An alternative approach is surrogate model assisted optimization [24, 25], which has particularly gained significant interest with the recent advances in machine learning [26, 27, 28, 29]. In this approach a surrogate model is trained on the PDE variables calculated with solutions from a PDE solver at several different parameter values. The surrogate model which is computationally cheaper to evaluate than the PDE solver, can then be used for optimization. However, this requires large amount of training data or PDE solves, and the model is only approximate. Another related approach is black box optimization, such as Bayesian optimization (see Section 4.2) where the PDE solver is treated as a black box and surrogate model construction and optimization are accomplished simultaneously by sequentially/adaptively sampling parameters for querying the PDE solver [30, 31, 32].
@@ 3.1.3 Optimization Problem
In Sec. 5 we propose BVQPCO, a bi-level variational quantum framework to solve the above optimization problem. The VQLS is used to solve the linear system constraint (57) for given parameters $l,\alpha$, and evaluate the ratio $\frac{\langle\mathbf{\phi}_{d},\tilde{\mathbf{u}}_{s}\rangle}{\langle\mathbf{\phi}_{n},\tilde{\mathbf{u}}_{s}\rangle}$ required for computing the cost function $C_{d}$, while a classical black box optimizer, e.g., Bayesian Optimization (BO) is used to select next set of parameters $l,\alpha$ values based on the evaluated cost. Since, VQLS itself involves optimization of the ansatz parameters which parameterizes the solution of the linear system, this leads to a sequence of nested bi-level optimization steps within a variational setting.
@@ 4 Methods
The two main building blocks of BVQPCO are VQLS and BO. In this section, we provide an overview of these concepts and discuss the specific choice of parameters and techniques for the heat transfer optimization problem discussed in Sec. 3.1.
@@ 4.1 Variational Quantum Linear Solver
where, $|\mathbf{b}\rangle$ denotes the quantum state prepared from the vector $\mathbf{b}$, and same for $|\mathbf{\psi}\rangle$. The inputs to this algorithm are the matrix $\mathbf{A}$ given as a linear combination of unitary (LCU) matrices $\mathbf{A}_{i}$ with the coefficients $\alpha_{i}$
@@ 4.1.1 Linear Combination of Unitaries Decomposition
with complex coefficients $\alpha_{i}$ as given in Eqn. 60. In this basis, each $\mathbf{A}_{i}\in P_{n}$ and its corresponding coefficient $\alpha_{i}$ can be determined via different approaches [34]. For instance, it can be shown that $\alpha_{i}=Tr(\mathbf{A}_{i}\cdot\mathbf{A})/2^{n}$. This matrix multiplication based approach can be computationally very expensive. A more efficient approach based on matrix slicing was proposed in [34] and we adapt it in this work. For the matrix in Eqn. 55, the number of terms scale almost linearly with the matrix size. The decomposition has 14 terms for a $8\times 8$ matrix, 26 terms for $16\times 16$ matrix and 54 terms for $32\times 32$ matrix.
@@ 4.1.2 State Preparation
General purpose methods exist to compute the unitary operator $\mathbf{U}_{b}$ [35, 36]. However, such methods may not always be efficient and can lead to unnecessarily deep circuits which in turn can affect the performance. Instead, we hand-craft circuits to encode states of form shown in  Eqn. 18 using only controlled rotations.  Fig. 2(a),(b) show the quantum circuits for the state on $4$ and $5$ qubits. The rotation angles are computed using user-defined values of the parameters in the state $\tilde{\mathbf{b}}$. In contrast, the quantum circuit for the same state constructed using Mottonen state preparation method is deeper as shown in Fig.  2 (c).
@@
Figure 2: State preparation with $T_{0}=2\big{(}\frac{\Delta t}{k\Delta x}q_{0}\big{)}$ using our method for (a) 3 qubits, $n_{x}=2,n_{t}=4$, and (b) 4 qubits $n_{x}=4,n_{t}=4$. (c) Mottonen state preparation using built-in PennyLane implementation for 4 qubits, $n_{x}=4,n_{t}=4$ case.

### v2

@@ Abstract
We present a novel variational quantum framework for linear partial differential equation (PDE) constrained optimization problems. Such problems arise in many scientific and engineering domains. For instance, in aerodynamics, the PDE constraints are the conservation laws such as momentum, mass and energy balance, the design variables are vehicle shape parameters and material properties, and the objective could be to minimize the effect of transient heat loads on the vehicle or to maximize the lift-to-drag ratio. The proposed framework utilizes the variational quantum linear system (VQLS) algorithm and a black box optimizer as its two main building blocks. VQLS is used to solve the linear system, arising from the discretization of the PDE constraints for given design parameters, and evaluate the design cost/objective function. The black box optimizer is used to select next set of parameter values based on this evaluated cost, leading to nested bi-level optimization structure within a hybrid classical-quantum setting. We present detailed computational error and complexity analysis to highlight the potential advantages of our proposed framework over classical techniques. We implement our framework using the PennyLane library, apply it to a heat transfer optimization problem, and present simulation results using Bayesian optimization as the black box optimizer.
@@ 1 Introduction
We present a novel variational quantum framework for optimization problems constrained by linear partial differential equations (PDE). Variational quantum algorithms (VQAs) are hybrid classical-quantum algorithms that have emerged as promising candidates to optimally utilize today’s Noisy Intermediate Scale Quantum (NISQ) devices. VQAs use a quantum subroutine to evaluate the cost/ energy function using a shallow parameterized quantum circuit. They provide a versatile framework to tackle a wide variety of problems. As such, they have been successfully used in many applications such as finding ground and excited states of Hamiltonians [1, 2, 3, 4], dynamical quantum simulation [5, 6], combinatorial optimization [7, 8], solving linear systems [9, 10, 11], integer factorization [12], principal component analysis [13, 14, 15] and quantum machine learning [16].
@@
In this work we are interested in the application of the variational quantum linear solver (VQLS) [9], a type of VQA for solving ordinary differential equations (ODEs) and partial differential equations (PDEs) [17, 18, 19, 20]. In this framework the ODEs/PDEs are discretized in space/time and embedded into a system of linear equations. VQLS is then applied to the linear system to obtain the normalized solution and compute observables of interest. While there are no theoretical guarantees associated with VQLS, its query complexity is empirically found to scale polylogarithmically with the size of the linear system [9] and thus could be beneficial over classical linear solvers that have a polynomial scaling. Unlike variational algorithms, fault tolerant quantum algorithms (FTQA) for ODE/PDE simulation have also been proposed and have provable theoretical guarantees/advantages over classical methods. Such FTQA rely either on quantum linear system solver or Hamiltonian simulation framework, see for example [21, 22, 23, 24, 25, 26, 27, 28, 29]. While, such FTQA can be used in our proposed PDE constrained optimization framework, in this paper we restrict to a variational setting with a focus on NISQ-era implementation. In particular, we extend the VQLS based ODE/PDE solution framework for linear PDE constrained design optimization problems.
Many science and engineering applications necessitate the solution of optimization problems constrained by physical laws that are described by systems of PDEs. Such problems arise in several domains including aerodynamics, computational fluid dynamics (CFD), material science, computer vision, and inverse problems [30]. For instance in aerodynamic design the PDE constraints are the conservation laws for momentum, mass, and energy, the design variables are aerodynamic shape parameters and material properties, and the objective is to minimize heat transfer or to maximize lift-to-drag ratio. Inverse problems on other hand typically involve calibrating PDE model parameters to match the given measurements. Closed-form solutions are generally unavailable for PDE constrained optimization problems, necessitating the development of numerical methods [31, 32, 30]. A variety of classical gradient-based and gradient-free numerical methods have been developed, and rely on repeated calls to the underlying PDE solver. Since PDE simulations are computationally expensive, using them within the design/optimization loop can become a bottleneck.
To address this challenge we propose a variational quantum framework where a quantum device is used for the computationally demanding PDE simulations and design/optimization is carried out on a classical computer. To best of our knowledge, this is the first work considering the simulation and design problem jointly in the context of quantum algorithms. The main contributions of this work are as follows:
@@
  • We present detailed computational error analysis for VQLS based solution of linear ODEs under explicit and implicit Euler discretization schemes. Furthermore by combining this error analysis with the empirically known results for query complexity of VQLS, we assess potential utility of our BVQPCO framework over the classical methods.
  • We implement the BVQPCO framework using the PennyLane library, apply it to solve a prototypical heat transfer optimization problem, and present simulation results using Bayesian optimization (BO) as the black box optimizer.
The paper is organized as follows. In Sec. 3 we motivate the general structure of PDE constrained design optimization problem through a detailed mathematical formulation of the 1D heat transfer optimization problem. We describe our BVQPCO framework and outline a pseudo-code for its implementation in Sec. 4.1. Sec. 4.2 and 4.3 provide an overview of VQLS and BO, respectively. We present the main results from the complexity analysis of the BVQPCO framework in Sec. 5; detailed proofs are provided in the Appendix. In  Sec. 6 we illustrate the BVQPCO framework on the heat transfer optimization problem and provide details on the implementation using the PennyLane library. We finally conclude in Sec. 7 with avenues for future work.
@@ 2 Notation
We will use standard braket notation, i.e. $|\mathbf{\psi}\rangle$ and $\langle\mathbf{\psi}|$ in representing the quantum states [33]. The inner product between two quantum states $|\mathbf{\psi}\rangle$ and $|\mathbf{\phi}\rangle$ will be denoted by $\langle\mathbf{\psi}|\mathbf{\phi}\rangle$. For a vector $\mathbf{x}$, we denote by $|\mathbf{x}\rangle=\frac{\mathbf{x}}{\|\mathbf{x}\|}$ as the corresponding quantum state. The trace norm $\rho(\mathbf{\psi},\mathbf{\phi})$ between two pure states $|\mathbf{\psi}\rangle$ and $|\mathbf{\phi}\rangle$ is defined as
@@ 3 PDE Constrained Optimization
One popular classical approach to solve PDE constrained optimization is via gradient based methods [31, 32, 30]. For example, the gradient of the cost function $C_{d}$ w.r.t to $\mathbf{p}$ can be obtained via finite differencing. However, that requires solving the PDE (8) at least $N_{p}+1$ times to approximate the derivative of the solution $\mathbf{u}$ with respect to the parameters. A more efficient approach is via adjoint formulation, where one needs to solve the PDE (8) only one time and an additional adjoint equation to compute the required gradient [31, 32]. However, given that this gradient based approach requires forward simulation(s) for each gradient step, it can be computationally intensive. An alternative approach is surrogate model assisted optimization [35, 36], which has particularly gained significant interest with the recent advances in machine learning [37, 38, 39, 40]. In this approach a surrogate model is trained on the PDE variables calculated with solutions from a PDE solver at several different parameter values. The surrogate model which is computationally cheaper to evaluate than the PDE solver, can then be used for optimization. However, this requires large amount of training data or PDE solves, and the model is only approximate. Another related approach is black box optimization, such as Bayesian optimization (see Section 4.3) where the PDE solver is treated as a black box and surrogate model construction and optimization are accomplished simultaneously by sequentially/adaptively sampling parameters for querying the PDE solver [41, 42, 43].
@@ 3.1.3 Optimization Problem
In Sec. 4.1 we propose BVQPCO, a bi-level variational quantum framework to solve the above optimization problem. The VQLS is used to solve the linear system constraint (57) for given parameters $l,\alpha$, and evaluate the ratio $\frac{\langle\mathbf{\phi}_{d},\tilde{\mathbf{u}}_{s}\rangle}{\langle\mathbf{\phi}_{n},\tilde{\mathbf{u}}_{s}\rangle}$ required for computing the cost function $C_{d}$, while a classical black box optimizer, e.g., Bayesian Optimization (BO) is used to select next set of parameters $l,\alpha$ values based on the evaluated cost. Since, VQLS itself involves optimization of the ansatz parameters which parameterizes the solution of the linear system, this leads to a sequence of nested bi-level optimization steps within a variational setting.
@@ 4 Methods
In this section we describe the general BVQPCO framework and provide an overview of VQLS and BO which are its two main building blocks. We also discuss the specific choice of parameters and methods within VQLS/BO for the heat transfer optimization problem discussed in the Sec. 3.1.
4.1 BVQPCO Framework
$g_{i}(\mathbf{p})\leq 0,i=1,\cdots,N_{e},$
where, $\mathbf{p}\in R^{N_{p}}$ are the design/optimization variables, $\mathbf{\psi}\in\mathbb{R}^{n_{s}}$ is the discretized solution vector of the underlying PDE represented in the form of the linear constraints (60), $g_{i}(\mathbf{p})$ are the constraints on $\mathbf{p}$, and $C_{d}$ is the cost function which can expressed as a function $f$ involving $\mathbf{\phi}_{d},\mathbf{\phi}_{n}\in\mathbb{R}^{n_{s}}$ as fixed vectors associated with the cost function.
Figure 2: Top: Schematic showing transformation of a PDE constrained optimization problem into a variational quantum form. Bottom: Flow diagram of the BVQPCO framework. VQLS uses an inner level optimization to solve the linear system constraints, arising from the discretization of the underlying PDEs, for given design parameters, and evaluate the quantities related to design cost/objective function. A black box optimizer, e.g., BO is used for the outer level optimization to select next set of parameters values based on the evaluated design cost.
2:Output: Optimal parameters: $\mathbf{p}^{*}$
3:Initialize $k=0$ and $\mathbf{p}^{0}=\mathbf{p}_{in}$.
5:while stopping criteria not met do
6:   Compute LCU $\{\alpha_{i}(\mathbf{p}^{k}),\mathbf{A}_{i}(\mathbf{p}^{k})\}_{i=1}^{n_{k}}$ of $\mathbf{A}(\mathbf{p}^{k})$, and an unitary $\mathbf{U}_{b}^{k}$ such that $|\mathbf{b}(\mathbf{p}^{k})\rangle=\mathbf{U}_{b}^{k}|\mathbf{0}\rangle$.
7:   Compute $\mathbf{\theta}_{*}^{k}=VQLS(\{\alpha_{i}(\mathbf{p}^{k}),\mathbf{A}_{i}(\mathbf{p}^{k})\}_{i=1}^{n_{k}},\mathbf{U}_{b}^{k},\mathbf{V}(\mathbf{\theta}),\gamma,n_{sh})$ using the Algorithm 2.
9:   Use classical black box optimizer, e.g. BO Algorithm 3, to select next $\mathbf{p}^{k+1}$ subject to constraints $g_{i}(\mathbf{p}^{k+1})\leq 0,i=0,\cdots,N_{e}$.
11:   $k\leftarrow k+1$
12:end while
13:Return $\mathbf{p}^{k}$
Algorithm 1 BVQPCO Algorithm.
To solve this problem in a variational quantum framework, we propose BVQPCO framework, where the outer optimization level iteratively selects $\mathbf{p}^{k}$ using a classical black box optimizer based on the cost function evaluated using solution $|\mathbf{\psi}\rangle$ obtained via VQLS for the linear system $\mathbf{A}(\mathbf{p}^{k})|\mathbf{\psi}\rangle=|\mathbf{b}(\mathbf{p}^{k})\rangle$. Thus, the optimization problem (58) can be expressed in a variational quantum form as
$g_{i}(\mathbf{p})\leq 0,i=1,\cdots,N_{e},$
where, $\mathbf{V}(\mathbf{\theta})$ is the VQLS ansatz (see Sec. 4.2) which parametrizes the solution of the linear system (60), and $|\mathbf{\phi}_{d}\rangle,|\mathbf{\phi}_{n}\rangle\in R^{n_{s}}$ define quantum analogues of the cost function parameters $\mathbf{\phi}_{d}$ and $\mathbf{\phi}_{n}$, respectively. Note that since,
the cost function (58) can be equivalently expressed in the quantum form (61). Fig. 2 shows the overall flow diagram of the BVQPCO framework. The steps to solve (61) are described in Algorithm 1. Some remarks follow:
  • VQLS in the line (7) is implemented as described in the Algorithm 2. The inputs include the linear combination of unitaries (LCU) decomposition of $\mathbf{A}(\mathbf{p})$ (see  Sec. 4.2.1), unitary $\mathbf{U}_{b}^{k}$ which prepares $|\mathbf{b}(\mathbf{p}^{k})\rangle$ (see Sec. 4.2.2), $\mathbf{V}(\mathbf{\theta})$ is the selected ansatz as discussed above (see Sec. 4.2.3), $\gamma$ is the stopping threshold (see  Sec. 4.2.4) and $n_{sh}$ is the number of shots used in VQLS cost function evaluation.
  • Depending on how $\mathbf{A}$ depends on the parameters $\mathbf{p}$, a parameter dependent LCU decomposition $\{\alpha_{i}(\mathbf{p}),\mathbf{A}_{i}\}$ can be pre computed once, thus saving computational effort. For instance, for the optimization problem (57), $\mathbf{A}$ for both the explicit (see Eqn. 36) and implicit (see Eqn. 55) discretization schemes has a separable structure. Thus, LCU for the matrices $\tilde{\mathbf{A}}_{f1},\tilde{\mathbf{A}}_{f2}$ (and similarly $\tilde{\mathbf{A}}_{b1},\tilde{\mathbf{A}}_{b2}$) can be computed once, and the LCU of $\mathbf{A}$ can be readily constructed for any choice of parameters $\mathbf{p}=(l,\alpha)$.
  • Note that VQLS provides solution of the given linear system only upto a normalization constant. However, due to the assumed structure of the cost function, it is invariant to the normalization constant as shown by the expression (64). In fact, there is no loss of generality in the assumed form, i.e., $\frac{\langle\mathbf{\phi}_{d},\mathbf{\psi}\rangle}{\langle\mathbf{\phi}_{n},\mathbf{\psi}\rangle}$, as one can always choose $\langle\mathbf{\phi}_{n},\mathbf{\psi}\rangle$ to depend only the initial condition which is known a priori as part of the problem specification.
  • For the outer optimization in line (9) one can use any black box optimization method [44]. We propose to use BO (see Sec. 4.3) which is a sequential design strategy for global optimization of black-box functions that does not require derivatives, is robust to noisy evaluation of the cost function, and uses exploration/exploitation trade off to find optimal solution with minimum number of function calls.
  • By using VQLS for solving the linear system associated with discretized PDE one can potentially obtain computational advantage over classical methods discussed in the Sec. 3. In the Sec. 5 we provide a detailed computational error analysis for VQLS based solution of linear ODEs. Since, there are no theoretical results available for query complexity of VQLS, we employ empirically known results along with our rigorous VQLS error analysis to assess potential advantage of the BVQPCO framework over classical methods.
  • Note that in the BVQPCO framework one can replace VQLS by a quantum linear system algorithm (QLSA) [45, 46, 47]. This could be beneficial since QLSA have provable exponential advantage over classical linear system solvers. However, since QLSA have high quantum resource needs, incorporating QLSA within our proposed framework may only be feasible in a fault tolerant quantum computing setting.
@@ 4.2 Variational Quantum Linear Solver
where, $|\mathbf{b}\rangle$ denotes the quantum state prepared from the vector $\mathbf{b}$, and same for $|\mathbf{\psi}\rangle$. The inputs to this algorithm are the matrix $\mathbf{A}$ given as a LCU matrices $\mathbf{A}_{i}$ with the coefficients $\alpha_{i}$
@@ 4.2.1 LCU Decomposition
with complex coefficients $\alpha_{i}$ as given in Eqn. 67. In this basis, each $\mathbf{A}_{i}\in P_{n}$ and its corresponding coefficient $\alpha_{i}$ can be determined via different approaches [49]. For instance, it can be shown that $\alpha_{i}=Tr(\mathbf{A}_{i}\cdot\mathbf{A})/2^{n}$. This matrix multiplication based approach can be computationally very expensive. A more efficient approach based on matrix slicing was proposed in [49] and we adapt it in this work. For the matrix in Eqn. 55, the number of terms scale almost linearly with the matrix size. The decomposition has 14 terms for a $8\times 8$ matrix, 26 terms for $16\times 16$ matrix and 54 terms for $32\times 32$ matrix.
@@ 4.2.2 State Preparation
General purpose methods exist to compute the unitary operator $\mathbf{U}_{b}$ [50, 51]. However, such methods may not always be efficient and can lead to unnecessarily deep circuits which in turn can affect the performance. Instead, we hand-craft circuits to encode states of form shown in  Eqn. 18 using only controlled rotations.  Fig. 3(a),(b) show the quantum circuits for the state on $3$ and $4$ qubits. The rotation angles are computed using user-defined values of the parameters in the state $\tilde{\mathbf{b}}$. In contrast, the quantum circuit for the same state constructed using Mottonen state preparation method is deeper as shown in Fig.  3 (c).
@@
Figure 3: Quantum circuit to prepare state $\tilde{\mathbf{b}}$ in Eqn. 18 with $T_{0}(x)=2\big{(}\frac{\Delta t}{k\Delta x}q^{1}\big{)},q^{i}=q^{1}=50\forall i,k=1.0,\Delta t=0.25,\Delta x=0.2$ (a) 3 qubits, $n_{x}=2,n_{t}=4$, and (b) 4 qubits $n_{x}=4,n_{t}=4$. (c) Mottonen state preparation using built-in PennyLane implementation for 4 qubits, $n_{x}=4,n_{t}=4$ case.